OpenStudy (anonymous):

A clock loses 32 minutes in a day. If the clock is set at 3 a.m. today, then what will be the correct time when the clock indicates 7 p.m. on the 4th day

OpenStudy (sirm3d):

minutes to recover : \[\frac{32}{1\cdot 24}=\frac{t}{4\cdot 24+16}\]

OpenStudy (anonymous):

Minutes to recover = 3*32 + 16/24 *32 = 117.333 minutes = 1 hour 57 minutes 20 seconds

OpenStudy (anonymous):

So, correct time will be 8 pm 57 minutes 20 seconds